24% of what is 104?

Found 2 solutions by richwmiller, solver91311:
Answer by richwmiller(17219)

.24x=104
24x=10400
x=10400/24=1300/3=433.333

Answer by solver91311(24713)



Let represent "what"

"of" means multiply, and

% means how many out of 100, so



Solve for

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism